rava:

Matt paid $55 at a restaurant. That amount included a 10% tip. What was the check amount before the tip? $40 $43 $45 $50 $52

Ultrilliam:

10% of 40 = 4, 10% of 50 = 5. 50+5=55, so 50$

Ultrilliam:

TIP, 10/40=4 10/50=5

Ultrilliam:

@rava There ya go

Ultrilliam:

@rava here